1.Find the value of 20x when x=10?
2.find the value of 3a when a=-4?
3.find the value of 11(a+b) when a=7 and b=3?


Answer by checkley71(8403)

1) 20X
20*10=200 ANSWER.
2) 3A
3*-4=-12 ANSWER.
3) 11(A+B)
11(7*3)=11*10=110 ANSWER.